# Periodic double tilings of the plane

### v2

@@ 3. Optimal lattice
Let $E$ be the hexagon obtained by prologing the six flat edges of $E_{2}$. This hexagon generates a periodic $1$-tiling ${\mathbf{T}}^{\prime}$ with the same lattice as ${\mathbf{T}}$. Notice, moreover, that $\operatorname{Per}({\mathbf{T}}^{\prime})$ is obtained from $\operatorname{Per}({\mathbf{T}})$ by subtracting the perimeter of the triangular tile $E_{1}$ and adding the length of a tripod on the vertices of $E_{1}$. At this point, we can repeat the proof of Lemma 3.1 because the length $\operatorname{Per}({\mathbf{T}})$ is given by the formula (3.1) with an additional term which only depends on $\left\lvert E_{1}\right\rvert$ and has no variation with respect to variations of the lattice.
@@
Let ${\mathbf{T}}$ be a non-degenerate, isoperimetric periodic $2$-tiling of $\mathbb{R}^{2}$. If ${\mathbf{T}}$ is a $(4;8)$ configuration, then the lattice admits a fundamental domain $D$ which is a square i.e., the lattice $G$ has two generators $g,g^{\prime}\in G$ with equal length the curvilinear rectangle, $E_{1}$ is a curvilinear square and the chipped octagon $E_{2}$ is a chipped square.
@@
Consider a system of coordinates, as in Figure 6, with the origin in the center of the curvilinear rectangle $E_{1}$, axes parallel to the axes of $E_{1}$ and such that the first coordinate is along the longest axis. In these coordinates the vertices of $E_{1}$ are (±$(\pm a,\pm b)$ for some $a>b>0$. The coordinates of the midpoints of the two flat edges emitted from the vertices $(a,\pm b)$ are $(u,v)$ and $(u^{\prime},v^{\prime})$ for some $u,v,u^{\prime},-v^{\prime}>0$ so that $g=2(u,v)$, $g^{\prime}=2(u^{\prime},v^{\prime})$ are two generators of the lattice $G$. Let $0<\theta\leq\frac{\pi}{12}$ be the angle between the arc and the chord of the edge with vertices $(a,\pm b)$. Clearly, by the 120-degrees condition, the lines containing the flat edges emitted from the vertices $(a,\pm b)$ define the same angle $\alpha$ with the $x$-axis, and an easy computation gives $\alpha=\frac{\pi}{6}+\theta$. Then:
@@
where $\ell$ is the distance
@@
$=8ab+4(a\sin\alpha+b\cos\alpha)(\ell+\ell^{\prime})+2\ell\ell^{\prime}\sin\alpha\cos\alpha$
-:$\eqcolon A(\ell,\ell^{\prime})-$
@@
Suppose by contradiction that $\ell\neq\ell^{\prime}$. If we take ¯$\bar{\ell}=\frac{\ell+\ell^{\prime}}{2}$ we have $P(\bar{\ell},\bar{\ell})=P(\ell,\ell^{\prime})$ and $A(\bar{\ell},\bar{\ell})>A(\ell,\ell^{\prime})$ because ¯2>superscript¯$\bar{\ell}^{2}>\ell\ell^{\prime}$. Hence there exists $\lambda<1$ such that $A(\lambda\bar{\ell},\lambda\bar{\ell})=A(\ell,\ell^{\prime})$ while $P(\lambda\bar{\ell},\lambda\bar{\ell})<P(\ell,\ell^{\prime})$. In this process $\left\lvert E_{1}\right\rvert$ is kept fixed, hence we have a contradiction with the minimality of $\operatorname{Per}({\mathbf{T}})$ with respect to a change of the lattice. The condition $\ell=\ell^{\prime}$ means that $D$ is a rhombus, as stated.
Step 2. We are going to prove that $a=b$ which means that $E_{1}$ is a curvilinear square, $E_{2}$ is a chipped square and the fundamental domain $D$ is a square. The configuration is determined by the parameters $a,b,u,v$ where $(a,b)$ determines the shape of $E_{1}$ and $(u,v)$ determines the lattice $G$. There is a compatibility condition between $a,b,u,v$ which is given by the fact that the segment from $(a,b)$ to $(u.v)$ must define equal angles of 120 degrees with the two circular edges of $E_{1}$ joining in $(a,b)$. We have:
@@
$\left\lvert E_{1}\right\rvert$ $=r^{2}\left(1-\sqrt{3}+\frac{\pi}{3}-2\sin^{2}\gamma\right)$
@@
$r$ $=r(a,b)=\sqrt{a^{2}+\sqrt{3}ab+b^{2}}.$
Let $m=\min\left\{\left\lvert E_{1}\right\rvert,\left\lvert D\right\rvert-\left\lvert E_{1}\right\rvert\right\}$ and let $E_{1}^{\prime}$ be the curvilinear square with angles of 120 degrees and vertices in the points (±$(\pm c,\pm c)$ with $c>0$ appropriately chosen so that $\left\lvert E_{1}^{\prime}\right\rvert=m$ i.e., by Lemma A.4,
$r^{2}(c,c)\left(1-\sqrt{3}+\frac{\pi}{3}\right)=c^{2}(2+\sqrt{3})\cdot\left(1-\sqrt{3}+\frac{\pi}{3}\right)=m.$
Take $w=\sqrt{uv}$ and consider the group $G^{\prime}$ generated by $2(w,w)$ and $2(w,-w)$ which has a fundamental domain $D^{\prime}$ which is a square with vertices $(0,0),2(w,w),2(w,-w),(4w,0)$ and the same area as $D$: $\left\lvert D^{\prime}\right\rvert=8w^{2}=8uv$. Notice that $\left\lvert E_{1}^{\prime}\right\rvert$ strictly contains a square of side $2c$, hence $\left\lvert E_{1}^{\prime\prime}\right\rvert>4c^{2}$ and since $\left\lvert E_{1}^{\prime}\right\rvert=m$ is smaller than half the area of $D$ we obtain that $c<w$. So, the translations $E_{1}^{\prime}+g^{\prime}$ with $g^{\prime}\in G^{\prime}$ are pairwise disjoint and we can define a the tile
@@
$P(x,z)=\frac{2}{3}\pi r(x,x)+2\sqrt{2}(z-x)=\left(\frac{2}{3}\pi\sqrt{2+\sqrt{3}}-2\sqrt{2}\right)x+2\sqrt{2}z$
and check that this function is increasing in $x$ so that $\operatorname{Per}({\mathbf{T}}^{\prime})=P(c,w)<P(d,w)$. Since $(d,d)=(a,b)$ are two points on the same level curve of $r(x,y)$ which is an ellipse $E$ while $(w,w)$ and $(u,v)$ are on the same level curve of the function $xy$ which is a hyperbola $H$, and clearly the smaller distance between $E$ and $H$ is obtained by taking the points on the axis $y=x$, we have that
$\sqrt{(u-a)^{2}+(v-b)^{2}}\geq\sqrt{2}\left\lvert d-w\right\rvert\geq d-w.$
@@
So we have $a=b$ and $u=v$ so that
$\operatorname{Per}({\mathbf{T}})$ $=\frac{2}{3}\pi\sqrt{2+\sqrt{3}}\cdot a+4\sqrt{2}(u-a)$
@@
$\operatorname{Per}({\mathbf{T}})$ $=\frac{2\pi-6\sqrt{4-2\sqrt{3}}}{\sqrt{9+4\pi-9\sqrt{3}}}\sqrt{\left\lvert E_{1}\right\rvert}+2\sqrt{\left\lvert D\right\rvert}$
@@
$m_{1}$ :-$\coloneq\sqrt{2}\sqrt{\pi-\sqrt{3}}:$ $q_{1}$ :-$4]{12}:$
[…]
@@
    if $x_{2}<x<\frac{1}{2}$ then ${\mathbf{T}}$ is a $(6;6)$ configuration and $\mathcal{I}(x)=q_{3}$;
@@ 4.2. Isoperimetric $N$-tilings
By [29, Theorem 5.2], we know that a periodic tiling composed by $N$ adjacent hexagons with almost the same amount of area is isoperimetric. Again, we stress that this was achieved by a qualitative argument and we do not know how close the areas should be for this to happen.
@@
Figure 7. The isoperimetric profile $\mathcal{I}(x)$ of the periodic $2$-tiling compared to the asymptotic profile of a non-periodic tiling made of hexagons with the same areas. See (4.1).
@@ 4.3. Kelvin problem with unequal cells
At least in the planar case, it is reasonable to ask if the partition $(E_{i})$ of $\mathbb{R}^{2}$ induced by an isoperimetric periodic tiling satisfying $\left\lvert E_{2i}\right\rvert=x$, $\left\lvert E_{2i+1}\right\rvert=1-x$ are local minimizers for the perimeter. It is easy to understand that a possible competitor $(F_{i}^{r})$ on a ball $B_{r}$ is obtained by subdividing the ball into two regions of measure proportional to $x$ and $1-x$ and fill each area with a honeycomb made of hexagons of area $x$ in the first region and area $1-x$ on the second one. One would have
@@
In Figure 7 we have superimposed the plot of $\mathcal{J}(x)$ with $\mathcal{I}(x)$. One can notice that $\mathcal{I}(x)=\mathcal{J}(x)$ only for $x\in\{0,x^{\prime},x^{\prime\prime},x^{\prime\prime\prime},1\}$ with $x^{\prime}\approx 0.038$, $x^{\prime\prime}\approx 0.105$ and $x^{\prime\prime\prime}\approx 0.171$ and $\mathcal{I}(x)\leq\mathcal{I}(x)$ for $x\in[0,x^{\prime}]\cup[x^{\prime\prime},x^{\prime\prime\prime}]$. Hence for all other values of $x$, the isoperimetric periodic tiling is not a local minimizer. This might suggest that minimizers exist only for some choices of the assigned areas, for example when we have the configurations $(3;9)$ and $(4;8)$. In particular, a natural question is understanding if our configurations are local minimizers in the range where they are better than the tiling with regular hexagons.
@@ Appendix A Useful geometric computations
Let $A=(0,0)$, $B=(u,0)$, and $C=(w,v)$ be three points in $\mathbb{R}^{2}$ and let $S$ be a point such that the three segment $AS$, $BS$, $CS$ define equal angles of 120 degrees in $S$. Then
@@
By Melzak’s construction the length of the tripod is equal to the distance $\left\lvert C-M\right\rvert$ where $M=\left(\frac{u}{2},-\frac{\sqrt{3}}{2}u\right)$ is the third vertex of the equilateral triangle with vertices in $A$ and $B$, opposite to $C$. The result follows. ∎
@@
$\operatorname{Per}(E)$ $=\frac{2}{3}\pi r.$
@@
Step 1. First of all, we claim that the quadrilateral must have the symmetries of a rectangle. If opposite arcs of the quadrilateral have different centers, then the claim easily follows from [41, Lemma 5.1] (just notice that if we fix an edge $e$ the two adjacent edges are contained in two circles which must be symmetric with respect to the axis of $e$). So it is enough to prove that opposite arcs cannot have the same center. Suppose by contradiction that they have the same center. Then the other pair of arcs must have the same center as well (simply because they have the same curvature radius, and the curvilinear quadrilateral is assumed to be convex). But then the region $E$ would be a disk, contradicting the assumption that the vertices have 120 degrees.
@@
By trigonometric formulas, the sought identity for $\left\lvert E\right\rvert$ follows. Finally, if $\theta=\frac{\pi}{12}$, then the inner rectangle is a square. ∎
@@
Acknowledgments. All authors are members of INDAM-GNAMPA. The authors acknowledge support from the MIUR Excellence Department Project awarded to the Department of Mathematics, University of Pisa, CUP I57G22000700001. M.N. acknowledges partial support from the PRIN 2022 project 2022E9CF89, and the PRIN 2022 PNRR project P2022WJW9H. E.P. acknowledges support from the PRIN 2022 project 2022PJ9EFL, and from the project PRA 2022 14 GeoDom (Università di Pisa).

### v3

@@ 3. Optimal lattice
Let $E$ be the hexagon obtained by prolonging the six flat edges of $E_{2}$. This hexagon generates a periodic $1$-tiling $\mathbf{T}$$^{\prime}$ with the same lattice as $\mathbf{T}$}. Notice, moreover, that $\operatorname{Per}({\mathbf{T}}^{\prime})$ is obtained from $\operatorname{Per}({\mathbf{T}})$ by subtracting the perimeter of the triangular tile $E_{1}$ and adding the length of a tripod on the vertices of $E_{1}$. At this point, we can repeat the proof of Theorem 3.1 because the length $\operatorname{Per}({\mathbf{T}})$ is given by the formula (3.1) with an additional term which only depends on $\left\lvert E_{1}\right\rvert$ and has no variation with respect to variations of the lattice.
@@
Let $\mathbf{T}$} be a non-degenerate, isoperimetric periodic $2$-tiling of $\mathbb{R}^{2}$. If $\mathbf{T}$} is a $(4;8)$ configuration, then the lattice admits a fundamental domain $D$ which is a square i.e., the lattice $G$ has two generators $g,g^{\prime}\in G$ with equal length the curvilinear rectangle, $E_{1}$ is a curvilinear square and the chipped parallelogram $E_{2}$ is a chipped square.
@@
Consider a system of coordinates, as in Figure 6, with the origin in the center of the curvilinear rectangle $E_{1}$, axes parallel to the axes of $E_{1}$ and such that the first coordinate is along the longest axis. In these coordinates the vertices of $E_{1}$ are $(\pm a,\pm b)$ for some $a>b>0$. The coordinates of the midpoints of the two flat edges emitted from the vertices $(a,\pm b)$ are $(u,v)$ and $(u^{\prime},v^{\prime})$ for some $u,v,u^{\prime},-v^{\prime}>0$ so that $g=2(u,v)$, $g^{\prime}=2(u^{\prime},v^{\prime})$ are two generators of the lattice $G$. Let $0<\theta\leq\frac{\pi}{12}$ be the angle between the arc and the chord of the edge with vertices $(a,\pm b)$, see Figure 8. Clearly, by the 120-degrees condition, the lines containing the flat edges emitted from the vertices $(a,\pm b)$ define the same angle $\alpha$ with the $x$-axis, and an easy computation gives $\alpha=\frac{\pi}{6}+\theta$. Let $r$ be the radius of curvature of the region $E_{1}$. Then:
@@
where $\ell$ is the distance from $A=(a,b)$ to $B=(u,v)$ and $l^{\prime}$ is the distance from (a,b)(a,-b) to $(u^{\prime},v^{\prime})$. Moreover
@@
$=8ab+4(a\sin\alpha+b\cos\alpha)(\ell+\ell^{\prime})+8\ell\ell^{\prime}\sin\alpha\cos\alpha$
$\eqcolon A(\ell,\ell^{\prime})$
@@
Suppose by contradiction that $\ell\neq\ell^{\prime}$. If we take $\bar{\ell}=\frac{\ell+\ell^{\prime}}{2}$ we have $P(\bar{\ell},\bar{\ell})=P(\ell,\ell^{\prime})$ and $A(\bar{\ell},\bar{\ell})>A(\ell,\ell^{\prime})$ because $\bar{\ell}^{2}>\ell\ell^{\prime}$. Hence there exists $\lambda<1$ such that $A(\lambda\bar{\ell},\lambda\bar{\ell})=A(\ell,\ell^{\prime})$ while $P(\lambda\bar{\ell},\lambda\bar{\ell})<P(\ell,\ell^{\prime})$. In this process $\left\lvert E_{1}\right\rvert$ is kept fixed, hence we have a contradiction with the minimality of $\operatorname{Per}({\mathbf{T}})$ with respect to a change of the lattice. The condition $\ell=\ell^{\prime}$ means that $D$ is a rhombus, as stated.
Step 2. We are going to prove that $a=b$ which means that $E_{1}$ is a curvilinear square, $E_{2}$ is a chipped square and the fundamental domain $D$ is a square. The configuration is determined by the parameters $a,b,u,v$ where $(a,b)$ determines the shape of $E_{1}$ and $(u,v)$ determines the lattice $G$. There is a compatibility condition between $a,b,u,v$ which is given by the fact that the segment from $(a,b)$ to $(u,v)$ must define equal angles of 120 degrees with the two circular edges of $E_{1}$ joining in $(a,b)$. By Lemma A.4 we have:
@@
$\left\lvert E_{1}\right\rvert$ $=r^{2}\left(1-\sqrt{3}+\frac{\pi}{3}-2\sin^{2}(\theta-\frac{\pi}{12})\right)$
@@
$r$ $=r(a,b)=2\sqrt{a^{2}+\sqrt{3}ab+b^{2}}.$
Let $m=\min\left\{\left\lvert E_{1}\right\rvert,\left\lvert D\right\rvert-\left\lvert E_{1}\right\rvert\right\}$ and let $E_{1}^{\prime}$ be the curvilinear square with angles of 120 degrees and vertices in the points $(\pm c,\pm c)$ with $c>0$ appropriately chosen so that $\left\lvert E_{1}^{\prime}\right\rvert=m$ i.e., by Lemma A.4,
$r^{2}(c,c)\left(1-\sqrt{3}+\frac{\pi}{3}\right)=4c^{2}(2+\sqrt{3})\cdot\left(1-\sqrt{3}+\frac{\pi}{3}\right)=m.$
Take $w=\sqrt{uv}$ and consider the group $G^{\prime}$ generated by $2(w,w)$ and 2(w,w)2(w,-w) which has a fundamental domain $D^{\prime}$ which is a square with vertices (0,0),2(w,w),2(w,w),(4w,0)(0,0),2(w,w),2(w,-w),(4w,0) and the same area as $D$: $\left\lvert D^{\prime}\right\rvert=8w^{2}=8uv$. Notice that $\left\lvert E_{1}^{\prime}\right\rvert$ strictly contains a square of side $2c$, hence $\left\lvert E_{1}^{\prime}\right\rvert>4c^{2}$ and since $\left\lvert E_{1}^{\prime}\right\rvert=m$ is smaller than half the area of $D$ we obtain that $c<w$. So, the translations $E_{1}^{\prime}+g^{\prime}$ with $g^{\prime}\in G^{\prime}$ are pairwise disjoint and we can define a the tile
@@
$P(x,z)=\frac{2}{3}\pi r(x,x)+4\sqrt{2}(z-x)=\left(\frac{4}{3}\pi\sqrt{2+\sqrt{3}}-4\sqrt{2}\right)x+4\sqrt{2}z$
and check that this function is increasing in $x$ so that $\operatorname{Per}({\mathbf{T}}^{\prime})=P(c,w)<P(d,w)$. Since $(d,d)$ and $(a,b)$ are two points on the same level curve of $r(x,y)$ which is an ellipse $E$ while $(w,w)$ and $(u,v)$ are on the same level curve of the function $xy$ which is a hyperbola $H$, and clearly the smaller signed distance between $E$ and $H$ is obtained by taking the points on the axis $y=x$, we have that
$\sqrt{(u-a)^{2}+(v-b)^{2}}\geq\sqrt{2}\left\lvert w-d\right\rvert\geq w-d.$
@@
So we have $a=b$ and $u=v$. Hence
$\operatorname{Per}({\mathbf{T}})$ $=P(a,u)=\left(\frac{4}{3}\pi\sqrt{2+\sqrt{3}}-4\sqrt{2}\right)\cdot a+4\sqrt{2}u$
@@
$\operatorname{Per}({\mathbf{T}})$ $=\frac{2\pi-6\sqrt{4-2\sqrt{3}}}{\sqrt{9+3\pi-9\sqrt{3}}}\sqrt{\left\lvert E_{1}\right\rvert}+2\sqrt{\left\lvert D\right\rvert}$
@@
$m_{1}$ $\coloneq\sqrt{2}\sqrt{\pi-\sqrt{3}}$ $q_{1}$ $\coloneq\sqrt[4]{12}$
$m_{2}$ $\coloneq 2\sqrt{\frac{\pi}{3}+1-\sqrt{3}}$ $q_{2}$ $\coloneq 2$
$q_{3}$ $\coloneq 2\sqrt[4]{3}$
@@
    if $x_{2}<x\leq\frac{1}{2}$ then $\mathbf{T}$} is a $(6;6)$ configuration and $\mathcal{I}(x)=q_{3}$;
@@ 4.2. Isoperimetric $N$-tilings
By [30, Theorem 5.2], we know that a periodic tiling composed by $N$ adjacent hexagons with almost the same amount of area is isoperimetric. Again, we stress that this was achieved by a qualitative argument and we do not know how close the areas should be for this to happen.
@@
Figure 7. Left: the isoperimetric profile $\mathcal{I}(x)$ of the periodic 22-tiling, compared with the asymptotic profile $\mathcal{J}(x)$, see (4.1), of a non-periodic tiling composed of hexagons with the same areas. Right: the difference $\mathcal{I}(x)-\mathcal{J}(x)$, highlighting the intervals where $\mathcal{I}(x)<\mathcal{J}(x)$.
@@ 4.3. Kelvin problem with unequal cells
At least in the planar case, it is reasonable to ask if the partition $(E_{i})$ of $\mathbb{R}^{2}$ induced by an isoperimetric periodic tiling satisfying $\left\lvert E_{2i}\right\rvert=x$, $\left\lvert E_{2i+1}\right\rvert=1-x$ are local minimizers for the perimeter. It is easy to understand that a possible competitor $(F_{i}^{r})$ on a ball $B_{r}$ is obtained by subdividing the ball into two regions of measure proportional to $x$ and 1x1-x and fill each area with a honeycomb made of hexagons of area $x$ in the first region and area 1x1-x on the second one. One would have
@@
In Figure 7 we have superimposed the plot of $\mathcal{J}(x)$ with $\mathcal{I}(x)$. One can notice that $\mathcal{I}(x)=\mathcal{J}(x)$ only for $x\in\{0,x^{\prime},x^{\prime\prime},x^{\prime\prime\prime},1\}$ with $x^{\prime}\approx 0.038$, $x^{\prime\prime}\approx 0.105$ and $x^{\prime\prime\prime}\approx 0.171$ and $\mathcal{I}(x)\leq\mathcal{J}(x)$ for $x\in[0,x^{\prime}]\cup[x^{\prime\prime},x^{\prime\prime\prime}]$. Hence for all other values of $x$, the isoperimetric periodic tiling is not a local minimizer. This might suggest that minimizers exist only for some choices of the assigned areas, for example when we have the configurations $(3;9)$ and $(4;8)$. In particular, a natural question is understanding if our configurations are local minimizers in the range where they are better than the tiling with regular hexagons.
@@ Appendix A Useful geometric computations
Let $A=(0,0)$, $B=(u,0)$, and $C=(w,v)$ be three points in $\mathbb{R}^{2}$ and let $S$ be a point such that the three segment $AS$, BSBS, CSCS define equal angles of 120 degrees in $S$. Then
@@
By Melzak’s construction [23] the length of the tripod is equal to the distance $\left\lvert C-M\right\rvert$ where $M=\left(\frac{u}{2},-\frac{\sqrt{3}}{2}u\right)$ is the third vertex of the equilateral triangle with vertices in $A$ and $B$, opposite to $C$. The result follows. ∎
@@
$\operatorname{Per}(E)$ $=\frac{2}{3}\pi r,$
$r$ $=2\sqrt{a^{2}+\sqrt{3}ab+b^{2}}.$
@@
Step 1. First of all, we claim that the quadrilateral must have the symmetries of a rectangle. If opposite arcs of the quadrilateral have different centers, then the claim easily follows from [42, Lemma 5.1] (just notice that if we fix an edge $e$ the two adjacent edges are contained in two circles which must be symmetric with respect to the axis of $e$). So it is enough to prove that opposite arcs cannot have the same center. Suppose by contradiction that they have the same center. Then the other pair of arcs must have the same center as well (simply because they have the same curvature radius, and the curvilinear quadrilateral is assumed to be convex). But then the region $E$ would be a disk, contradicting the assumption that the vertices have 120 degrees.
@@
By trigonometric formulas, the sought identity for $\left\lvert E\right\rvert$ follows. Finally, if $\theta=\frac{\pi}{12}$, then the inner rectangle is a square.
Step 3. Let us compute the curvature radius $r$ and conclude the proof. We claim that the centre $P$ of the arc passing through the points $A=(a,b)$ and (a,b)(a,-b) has coordinates $P=(-a-\sqrt{3}b,0)$, while the center $Q$ of the arc passing through $(a,b)$ and (a,b)(a,-b) has coordinates $Q=(0,-b-\sqrt{3}a)$. In fact it is a nice exercise in planar geometry to prove that if we construct two equilateral triangles on the exterior of two adjacent edges of a straight rectangle, the two constructed vertices ($P$ and $Q$ in our case) together with the opposite vertex of the rectangle ($A=(a,b)$ in our case), form an equilateral triangle. This means that the arcs centred in these two points and passing through the point $A$ have the same radius and define an angle of 120 degrees in $A$. It is not difficult to show that the points $P$, $Q$ are uniquely defined by these properties, obtaining the claim. Hence we have
$r^{2}=\left\lvert A-P\right\rvert^{2}=(a+a+\sqrt{3}b)^{2}+(b-0)^{2}=4a^{2}+3b^{2}+4\sqrt{3}ab+b^{2}=4a^{2}+4\sqrt{3}ab+4b^{2}.$
@@
Acknowledgments. All authors are members of INDAM-GNAMPA. The authors acknowledge support from the MIUR Excellence Department Project awarded to the Department of Mathematics, University of Pisa, CUP I57G22000700001. M.N. acknowledges partial support from the PRIN 2022 project 2022E9CF89, and the PRIN 2022 PNRR project P2022WJW9H. E.P. acknowledges support from the PRIN 2022 project 2022PJ9EFL, and from the project PRA 2022 14 GeoDom (Università di Pisa). The authors thank the anonymous referee for valuable comments that significantly improved the manuscript.